OpenStudy (anonymous):

2x^2+12x+k k>0 Find all values of "k" so that the trinomial can be factored as two binomials using integers? show work.

OpenStudy (anonymous):

ok break up 12 into as many sums as u can

OpenStudy (anonymous):

0+12 1+11 2+10 and so on

OpenStudy (amistre64):

1,12;2,6;3,4

OpenStudy (amistre64):

ack!!...wrong thought lol

OpenStudy (amistre64):

x+y = 12 looks like the equation of a line...

OpenStudy (anonymous):

suppose ur pair is a+b, then if ab is even, u can evaluate k=ab/2

OpenStudy (anonymous):

get it

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

just break up 12 into pairs of sums

OpenStudy (anonymous):

like 0 + 12 1+11 2+10 and so on

OpenStudy (anonymous):

and then multiply the 2 terms in each of the pairs and divide them by 2 to get k

OpenStudy (anonymous):

i still dont get it.....

OpenStudy (anonymous):

c wat do u do while factoring a quadratic eqn like this

OpenStudy (anonymous):

lets say its x^2 + 5x +6

OpenStudy (anonymous):

we dont use the quadratic equation at all until my junior year im only a freshmen

OpenStudy (anonymous):

hey this is called a quadratic equation..u just dunno the name

OpenStudy (anonymous):

my teacher doesnt want us useing quadratic equation to solve these

OpenStudy (anonymous):

so when u factor this expression out...u first multiply the c oefficient of the first term with the third

OpenStudy (anonymous):

u just lisn 2 me..were not using em

OpenStudy (anonymous):

i am listening sorry im confused

OpenStudy (anonymous):

no prob..

OpenStudy (anonymous):

when u factor this expression out...u first multiply the c oefficient of the first term with the third..and then u break the middle term into a sum, for eg if its x^2 + 5x + 6 the product is 1x6=6 so we break 5x into 2x + 3x, see 2.3 is 6 the same as 1.6 , dont we do that?

OpenStudy (anonymous):

u get it??like i did so many yesterday

OpenStudy (anonymous):

no i dont get it

OpenStudy (anonymous):

then m really sorry..i dont know how tto explain it..it really breaks my heart..bt ill giv u some advice

OpenStudy (anonymous):

try doing the factoring out questions urself...like i did ystrde..ul learn a lot..try noticing how ur factoring..

OpenStudy (anonymous):

i dont even know how to start htem so kinda hard to do something when your teacher that gets paid to teach doesnt teach at all

OpenStudy (anonymous):

shall i teach u?

OpenStudy (anonymous):

frm d start?

OpenStudy (anonymous):

you dont know any other way to explain this and im still confused